OpenStudy (anonymous):

A bag contains 6 cherry, 3 orange, and 2 lemon candies. You reach in and take 3 pieces of candy at random. The probability that you have 1 cherry candy and 2 lemon candies is

OpenStudy (anonymous):

This is a bit more difficult than usual because it is sampling without replacement. Say you pick 1 lemon on the first try, then there is only 1 lemon left and also the total changes, so the probabilities change. order: CLL P = 6/11 times 2/10 times 1/9 = 12 / 11 x 10 x 9 order: LCL P = 2/11 times 6/10 times 1/9 = 12 / 11 x 10 x 9 order: LLC P = 2/11 times 1/10 times 6/9 = 12 / 11 x 10 x 9 We have (12 + 12 + 12) / 11 X 10 X 9 = 36/990